# Neural Conditional Probability for Inference

### v1

@@ 6 Experiments
Conditional density estimation   We applied our NCP method to a benchmark of several conditional density models including those of Rothfuss et al., (2019); Gao and Hastie, (2022). See App. C.1 for the complete description of the data models and the complete list of compared methods in Tab. 2 with references. We also plotted several conditional CDF along with our NCP estimators in Fig. 4. To assess the performance of each method, we use Kolmogorov-Smirnov (KS) distance between the estimated and the true conditional CDFs. We test each method on nineteen different conditional values uniformly sampled between the 5%- and 95%-percentile of $p(x)$ and computed the averaged performance over all the used conditioning values. In Tab. 1, we report mean performance (KS distance $\pm$ std) computed over 10 repetitions, each with a different seed. NCP with whitening (NCP–W) outperforms all other methods on 4 datasets, ties with FlexCode (FC) on 1 dataset, and ranks a close second on another one behind NF. These experiments underscore NCP’s consistent performance. We also refer to Tab. 3 in App C.1 for an ablation study on post-treatments for NCP.
@@
Figure 1: Top: Laplace distribution; Bottom: Cauchy distribution. Conditional mean (top only) and $90\%$ confidence interval for NCP, NFs and CCP.
@@
In Fig. 1, observe first that the CCP regression achieves the best estimation of the conditional mean $mse=3.6\cdot 10^{-3}$ against $mse=3.8\cdot 10^{-2}$ for NFs and $mse=8.3\cdot 10^{-3}$ for NCP, as expected since the CCP regression model is well-specified in this example. However, the CCP confidence intervals are unreliable for most of the considered conditioning. We also notice instability for NF and CCP when conditioning in the neighborhood of $x=0$, with the NF confidence region exploding at $x=0$. We suspect this is due to the fact that the conditional distribution at $x=0$ is degenerate, hence violating the condition of existence of a diffeomorphism with the generating prior, a fundamental requirement for NFs models to work at all. Comparatively, NCP does not exhibit such instability around $x=0$; it only tends to overestimate the confidence region for conditioning close to $x=0$. The Cauchy distribution is known to be more challenging due to its heavy tail and undefined moments. In Fig 1 (bottom), we notice that NF and CCP completely collapse. This is not a surprising outcome since CCP relies on estimation of the mean which is undefined in this case, creating instability in the constructed confidence regions, while NF attempts to build a diffeomorphism between a Gaussian prior and the final Cauchy distribution. We suspect the conservative confidence region produced by NF might originate from the successive Jacobians involved in the NF mapping taking large values. In comparison, our NCP method still returns some reasonable results. Although the NCP coverage might appear underestimated for larger $x$, actual mean coverages computed on a test set of 200 samples are 88% for NCP, 99% for NF and 79% for CCP.
Finally, Tab. 5 in App. C.2 provides a comparison study on real data for learning a confidence region with NCP, NF and a split conformal predictor featuring a Random Forest regressor (RFSCP).
@@ 7 Conclusion
We introduced NCP, a novel neural operator approach to learn the conditional probability distribution from complex and highly nonlinear data. Our learning framework offers a number of benefits. Notably, it streamlines the training process by requiring just one unconditional training phase to learn the joint distribution $p(x,y)$. Subsequently, it allows us to efficiently derive conditional probabilities and other relevant statistics from the trained model analytically, without any additional conditional training steps or Monte Carlo sampling. Additionally, our method is backed by theoretical non-asymptotic guarantees ensuring the soundness of our training method and the accuracy of the obtained conditional statistics.
Our experiments on learning conditional densities and confidence regions demonstrate our approach’s superiority or equivalence to leading methods, even using a simple Multi-Layer Perceptron (MLP) with two hidden layers and GELU activations. This highlights the effectiveness of a minimalistic architecture coupled with a theoretically grounded loss function. Our results showcase the versatility of our approach as it can achieve competitive results without compromising performance over a wide variety of datasets and models. However the price we pay for this generality appears to be the need for a relatively large sample size ($n\gtrsim 10^{4}$) to start outperforming other methods. Hence, a future direction is to study how to incorporate prior knowledge into our method to make it more data-efficient. Future works will also investigate the performance of NCP for multi-dimensional time series, causality and more general sensitivity analysis in uncertainty quantification.
@@ Appendix B Proofs of theoretical results
B.1 Proof of Lemma 1
@@
$\mathbb{P}[Y\in B\,|\,X\in A]-\mathbb{P}[Y\in B]-\frac{{\langle\mathbb{1}_{A},[\![{\mathsf{D}_{Y|X}}]\!]_{r}\mathbb{1}_{B}\rangle}}{\mathbb{P}[X\in A]}=\frac{{\langle\mathbb{1}_{A},({\mathsf{D}_{Y|X}}-[\![{\mathsf{D}_{Y|X}}]\!]_{r})\mathbb{1}_{B}\rangle}}{\mathbb{P}[X\in A]}.$
@@
$|{\langle\mathbb{1}_{A},({\mathsf{D}_{Y|X}}-[\![{\mathsf{D}_{Y|X}}]\!]_{r})\mathbb{1}_{B}\rangle}|$ $\leq\lVert{\mathsf{D}_{Y|X}}-[\![{\mathsf{D}_{Y|X}}]\!]_{r}\rVert_{L^{2}_{\nu}(\mathcal{Y})\rightarrow L^{2}_{\mu}(\mathcal{X})}\lVert\mathbb{1}_{A}\rVert_{L^{2}_{\mu}(\mathcal{X})}\lVert\mathbb{1}_{B}\rVert_{L^{2}_{\nu}(\mathcal{Y})}$
$=\lVert{\mathsf{D}_{Y|X}}-[\![{\mathsf{D}_{Y|X}}]\!]_{r}\rVert_{L^{2}_{\nu}(\mathcal{Y})\rightarrow L^{2}_{\mu}(\mathcal{X})}\sqrt{\mathbb{P}[X\in A]}\sqrt{\mathbb{P}[Y\in B]},$
where the operator norm $\lVert{\mathsf{D}_{Y|X}}-[\![{\mathsf{D}_{Y|X}}]\!]_{r}\rVert_{L^{2}_{\nu}(\mathcal{Y})\rightarrow L^{2}_{\mu}(\mathcal{X})}$ is upper bounded by $\sigma^{\star}_{r+1}$ by definition of the SVD of ${\mathsf{D}_{Y|X}}$. ∎
B.2 Proof of Theorem 1
@@
In the following, to simplify notation, whenever dependency on the parameters is not crucial, recalling that $(X,Y)$ and $(X^{\prime},Y^{\prime})$ are two iid samples from the joint distribution $\rho$, we will denote the vector-valued random variables in the latent (embedding) space as $u:=u^{\theta}(X)$, $u^{\prime}:=u^{\theta}(X^{\prime})$, $v:=v^{\theta}(Y)$ and $v^{\prime}:=v^{\theta}(Y^{\prime})$, as well as $s=\sigma^{\theta}$ and $S:=S_{\theta}$. Then, we can write the training loss simply as $\mathbb{E}\,[L_{\gamma}(u,u^{\prime},v,v^{\prime},S)]$.
First, let us prove that $\mathcal{L}_{0}(\theta)=\lVert U_{\theta}S_{\theta}V_{\theta}^{*}\rVert_{\rm{HS}}^{2}-2\operatorname{\text{\rm tr}}(S_{\theta}U_{\theta}^{*}{\mathsf{D}_{Y|X}}V_{\theta})$. Indeed, we have that
$U_{\theta}^{*}{\mathsf{D}_{Y|X}}V_{\theta}=U_{\theta}^{*}{\mathsf{E}_{Y|X}}V_{\theta}-U_{\theta}^{*}\mathbb{1}_{\mathcal{X}}\otimes(V_{\theta}^{*}\mathbb{1}_{\mathcal{Y}})=\mathbb{E}[u^{\theta}(X)\mathbb{E}[v^{\theta}(Y)^{\top}\,|\,X]]-(\mathbb{E}[u^{\theta}(X)])(\mathbb{E}[v^{\theta}(Y)])^{\top},$
that is $U_{\theta}^{*}{\mathsf{D}_{Y|X}}V_{\theta}=\mathbb{E}[uv^{\top}]-\mathbb{E}[u]]\mathbb{E}[v]^{\top}$ is simply centered cross-covariance in the embedding space. Recalling that $U_{\theta}^{*}U_{\theta}=\mathbb{E}[uu^{\top}]$ and $V_{\theta}^{*}V_{\theta}=\mathbb{E}[vv^{\top}]$ are covariance matrices in the embedding space, we have that
$\mathcal{L}_{0}(\theta)$ $=-2\operatorname{\text{\rm tr}}\mathbb{E}[(S^{1/2}u)(S^{1/2}v)^{\top}]+2\operatorname{\text{\rm tr}}(\mathbb{E}[S^{1/2}u]\,\mathbb{E}[S^{1/2}v]^{\top})$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+\operatorname{\text{\rm tr}}(\mathbb{E}[(S^{1/2}u)(S^{1/2}u)^{\top}]\mathbb{E}[(S^{1/2}v)(S^{1/2}v)^{\top}])$
$=-2\mathbb{E}[u^{\top}Sv]+2(\mathbb{E}[u])^{\top}S(\mathbb{E}[v])+\operatorname{\text{\rm tr}}(\mathbb{E}[(S^{1/2}u)(S^{1/2}u)^{\top}]\mathbb{E}[(S^{1/2}v)(S^{1/2}v)^{\top}])$
@@
$\mathbb{E}\left[-uSv-u^{\prime}Sv^{\prime}+u^{\prime}Sv+uSv^{\prime}+\tfrac{1}{2}\operatorname{\text{\rm tr}}\left(S^{1/2}uu^{\top}Sv^{\prime}v^{\prime\top}S^{1/2}+S^{1/2}u^{\prime}u^{\prime\top}Svv^{\top}S^{1/2}\right)\right]$
$=\mathbb{E}\left[\tfrac{1}{2}\left(u^{\top}Sv^{\prime}\right)^{2}+\tfrac{1}{2}\left(u^{\prime\top}Sv\right)^{2}-(u-u^{\prime})S(v-v^{\prime})\right]$
$=\mathbb{E}\,[L_{0}(u,u^{\prime},v,v^{\prime},s)]=\mathbb{E}\,[L_{0}(u^{\theta}(X),u^{\theta}(X^{\prime}),v^{\theta}(Y),v^{\theta}(Y^{\prime}),\sigma^{\theta})].$
which implies that $\mathcal{L}_{0}(\theta)=\lVert U_{\theta}S_{\theta}V_{\theta}^{*}\rVert_{\rm{HS}}^{2}-2\operatorname{\text{\rm tr}}(S_{\theta}U_{\theta}^{*}{\mathsf{D}_{Y|X}}V_{\theta})$. Moreover, to show that $\mathcal{L}_{\gamma}(\theta)=\mathcal{L}_{0}(\theta)+\gamma\,\mathcal{R}(\theta)$.
It suffices to note that
@@
$\mathcal{L}_{0}^{r}(\theta):=\left\|D_{r}\!-\!U_{\theta}S_{\theta}V_{\theta}\right\|^{2}_{\rm HS}-\left\|D_{r}\right\|^{2}_{\rm HS}.$ (22)
@@
$\mathcal{L}_{0}^{r}(\theta)=-2\operatorname{\text{\rm tr}}\left(S_{\theta}^{1/2}U_{\theta}^{*}D_{r}V_{\theta}S_{\theta}^{1/2}\right)+\left\|U_{\theta}S_{\theta}V_{\theta}^{*}\right\|^{2}_{\rm HS},$
@@
$\mathcal{L}_{0}^{r}(\theta)=\lVert U_{\theta}S_{\theta}V_{\theta}^{*}\rVert_{\rm{HS}}^{2}-2\operatorname{\text{\rm tr}}(S_{\theta}^{1/2}U_{\theta}^{*}D_{r}V_{\theta}S_{\theta}^{1/2})=\mathcal{L}_{0}(\theta)+2\operatorname{\text{\rm tr}}(S_{\theta}U_{\theta}^{*}({\mathsf{D}_{Y|X}}-D_{r})V_{\theta}.$
@@
$\lvert\mathcal{L}_{0}^{r}(\theta)-\mathcal{L}_{0}(\theta)\rvert\leq\lvert\operatorname{\text{\rm tr}}(S_{\theta}U_{\theta}^{*}({\mathsf{D}_{Y|X}}\!-\!D_{r})V_{\theta})\rvert\leq\lVert S_{\theta}\rVert\lVert U_{\theta}^{*}\rVert_{\rm HS}\lVert{\mathsf{D}_{Y|X}}-[\![{\mathsf{D}_{Y|X}}]\!]_{r}\rVert\lVert V_{\theta}^{*}\rVert_{\rm HS},$

### v2

@@ 6 Experiments
Conditional density estimation   We applied our NCP method to a benchmark of several conditional density models including those of Rothfuss et al., (2019); Gao and Hastie, (2022). See Appendix C.1 for the complete description of the data models and the complete list of compared methods in Tab. 2 with references. We also plotted several conditional CDF along with our NCP estimators in Fig. 6. To assess the performance of each method, we use Kolmogorov-Smirnov (KS) distance between the estimated and the true conditional CDFs. We test each method on nineteen different conditional values uniformly sampled between the 5%- and 95%-percentile of $p(x)$ and computed the averaged performance over all the used conditioning values. In Tab. 1, we report mean performance (KS distance $\pm$ std) computed over 10 repetitions, each with a different seed. NCP with whitening (NCP–W) outperforms all other methods on 4 datasets, ties with FlexCode (FC) on 1 dataset, and ranks a close second on another one behind NF. These experiments underscore NCP’s consistent performance. We also refer to Tab. 3 in App C.1 for an ablation study on post-treatments for NCP.
@@
Figure 1: Conditional mean (top only) and 90% confidence interval for NCP, NFs and CCP. Top: Laplace distribution; Bottom: Cauchy distribution.
@@
In Fig. 1, observe first that the CCP regression achieves the best estimation of the conditional mean $mse=3.6\cdot 10^{-3}$ against $mse=3.8\cdot 10^{-2}$ for NFs and $mse=8.3\cdot 10^{-3}$ for NCP, as expected since the CCP regression model is well-specified in this example. However, the CCP confidence intervals are unreliable for most of the considered conditioning. We also notice instability for NF and CCP when conditioning in the neighborhood of $x=0$, with the NF confidence region exploding at $x=0$. We suspect this is due to the fact that the conditional distribution at $x=0$ is degenerate, hence violating the condition of existence of a diffeomorphism with the generating prior, a fundamental requirement for NFs models to work at all. Comparatively, NCP does not exhibit such instability around $x=0$; it only tends to overestimate the confidence region for conditioning close to $x=0$. The Cauchy distribution is known to be more challenging due to its heavy tail and undefined moments. In Fig 1 (bottom), we notice that NF and CCP completely collapse. This is not a surprising outcome since CCP relies on estimation of the mean which is undefined in this case, creating instability in the constructed confidence regions, while NF attempts to build a diffeomorphism between a Gaussian prior and the final Cauchy distribution. We suspect the conservative confidence region produced by NF might originate from the successive Jacobians involved in the NF mapping taking large values. In comparison, our NCP method still returns some reasonable results. Although the NCP coverage might appear underestimated for larger $x$, actual mean coverages computed on a test set of 200 samples are 88% for NCP, 99% for NF and 79% for CCP. Tab. 5 in Appendix C.2 provides a comparison study on real data for learning a confidence region with NCP, NF and a split conformal predictor featuring a Random Forest regressor (RFSCP).
[…]
High-dimensional synthetic experiment   We simulated the following $d$-distribution for different values of $d\in\{100,500,1000\}$. Let $\bar{x}=(\bar{x}_{1},\bar{x}_{2},0,\ldots,0)^{\top}\in\mathbb{R}^{d}$ where $x^{\prime}=(\bar{x}_{1},\bar{x}_{2})$ admits uniform distribution on the 2-dimensional unit sphere. We pick a random mapping $A\in\mathcal{O}_{d}$ and we set $X=A\bar{x}$ and the angle $\theta(X)=\arcsin(\bar{x}_{2})$. Next we consider two conditional distribution models for $Y|X$ (Gaussian and discrete) described in Figure 3. NCP performs similarly to NF in the Gaussian case and outperforms NF for discrete distribution. Figure 7 in Appendix C.3 demonstrates that NCP scales effectively with increasing dimensionality $d$. As the dimension rises from $d=100$ to $d=1000$, the computation time increases by only 20%, while maintaining strong statistical performance throughout.
Figure 3: High-dimensional synthetic experiment. We consider two models for $Y|X$ with $d=100$. Left: $Y|X\sim N(\theta(X),\sin(\theta(X)/2)$. Right: $Y\in\{1,2,3,4,5\}$ admits discrete distribution depending on $\theta(X)$: $Y|X\sim P_{1}$ if $\theta(X)\in[0,\pi/2)$, $P_{2}$ if $\theta(X)\in[\pi/2,\pi)$, $P_{3}$ if $\theta(X)\in[\pi,3\pi/2)$, $P_{4}$ if $\theta(X)\in[3\pi/2,2\pi)$. We take $P_{1}=(1/5,1/5,1/5,1/5,1/5)$, $P_{2}=(1/2,1/2,0,0,0)$, $P_{3}=(0,0,1,0,0)$, $P_{4}=(0,0,0,1/2,1/2)$.
High-dimensional experiment in molecular dynamics   We investigate protein folding dynamics and predict conditional transition probabilities between metastable states. Figure 2 shows how, by integrating our NCP approach with a graph neural network (GNN), we achieve accurate state forecasting and strong uncertainty quantification, enabling efficient tracking of transitions. For further context and a full model description, see App C.3.
@@ 7 Conclusion
We introduced NCP, a novel neural operator approach to learn the conditional probability distribution from complex and highly nonlinear data. NCP offers a number of benefits. Notably, it streamlines the training process by requiring just one unconditional training phase to learn the joint distribution $p(x,y)$. Subsequently, it allows us to efficiently derive conditional probabilities and other relevant statistics from the trained model analytically, without any additional conditional training steps or Monte Carlo sampling. Additionally, our method is backed by theoretical non-asymptotic guarantees ensuring the soundness of our training method and the accuracy of the obtained conditional statistics. Our experiments on learning conditional densities and confidence regions demonstrate our approach’s superiority or equivalence to leading methods, even using a simple Multi-Layer Perceptron (MLP) with two hidden layers and GELU activations. This highlights the effectiveness of a minimalistic architecture coupled with a theoretically grounded loss function. While complex architectures often dominate advanced machine learning, our results show that simplicity can achieve competitive results without compromising performance. Our numerical experiments suggest that, while our approach works well across different datasets and models, the price we pay for this generality appears to be the need for a relatively large sample size ($n\gtrsim 10^{4}$) to start outperforming other methods. Hence, a future direction is to study how to incorporate prior knowledge into our method to make it more data-efficient. Future works will also investigate the performance of NCP for multi-dimensional time series, causality and more general sensitivity analysis in uncertainty quantification.
Acknowledgements
We acknowledge financial support from EU Project ELIAS under grant agreement No. 101120237, by NextGenerationEU and MUR PNRR project PE0000013 CUP J53C22003010006 “Future Artificial Intelligence Research (FAIR)” and by NextGenerationEU and MUR PNRR project RAISE “Robotics and AI for Socio-economic Empowerment” (ECS00000035).
@@ Appendix B Proofs of theoretical results
B.1 A reminder on Hilbert spaces and compact operators
Definition 1.
[…]
 is complete with respect to the norm $\mathcal{H}\text{ is complete with respect to the norm }\|x\|=\sqrt{\langle x,x\rangle}\text{ for all }x\in\mathcal{H}.$
An important example of an infinite-dimensional Hilbert space is $L^{2}_{\mu}(\mathbb{R})$, the space of square-integrable functions w.r.t probability measure $\mu$ on $\mathbb{R}$ with the inner product defined as $\langle f,g\rangle=\int_{\mathbb{R}}f(x)\overline{g(x)}\,\mu(dx).$
Definition 2 (Bounded Operators).
Let $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ be Hilbert spaces. A linear operator $T:\mathcal{H}_{1}\to\mathcal{H}_{2}$ is called bounded if there exists a constant $C\geq 0$ such that for all $x\in\mathcal{H}_{1}$, the following inequality holds:
$\|Tx\|_{\mathcal{H}_{2}}\leq C\|x\|_{\mathcal{H}_{1}}.$
The smallest such constant $C$ is called the operator norm of $T$, denoted by $\|T\|$, and is given by:
$\|T\|=\sup_{x\neq 0}\frac{\|Tx\|_{\mathcal{H}_{2}}}{\|x\|_{\mathcal{H}_{1}}}.$
Bounded operators are continuous and play a key role in functional analysis.
Definition 3 (Compact Operators).
Let $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ be Hilbert spaces. A bounded linear operator $T:\mathcal{H}_{1}\to\mathcal{H}_{2}$ is called compact if for any bounded sequence $\{x_{n}\}\subset\mathcal{H}_{1}$, there exists a subsequence $\{x_{n_{k}}\}$ such that $Tx_{n_{k}}$ converges in $\mathcal{H}_{2}$.
Compact operators can be viewed as infinite-dimensional analogues of matrices with finite rank in finite-dimensional spaces.
A key result in the theory of compact operators is the existence of a singular value decomposition (SVD) for compact operators. The following is the statement of the Eckart-Young-Mirsky theorem:
Theorem 3 (Eckart-Young-Mirsky).
Let $T:\mathcal{H}_{1}\to\mathcal{H}_{2}$ be a compact operator between Hilbert spaces. Then $T$ can be decomposed as:
$T=\sum_{i=1}^{\infty}\sigma_{i}\langle\cdot,u_{i}\rangle v_{i},$
where $\{u_{i}\}\subset\mathcal{H}_{1}$ and $\{v_{i}\}\subset\mathcal{H}_{2}$ are orthonormal sets, and $\sigma_{i}$ are the singular values of $T$, which satisfy $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq 0$.
Moreover, for any rank-$k$ operator $T_{k}=\sum_{i=1}^{k}\sigma_{i}\langle\cdot,u_{i}\rangle v_{i}$, we have:
$\|T-T_{k}\|=\min_{\text{rank}(S)\leq k}\|T-S\|,$
[…]
B.2 Proof of Lemma 1
@@
$\mathbb{P}[Y\in B\,|\,X\in A]-\mathbb{P}[Y\in B]-\frac{{\langle\mathbb{1}_{A},[\![{\mathsf{D}_{Y|X}}]\!]_{d}\mathbb{1}_{B}\rangle}}{\mathbb{P}[X\in A]}=\frac{{\langle\mathbb{1}_{A},({\mathsf{D}_{Y|X}}-[\![{\mathsf{D}_{Y|X}}]\!]_{d})\mathbb{1}_{B}\rangle}}{\mathbb{P}[X\in A]}.$
@@
$|{\langle\mathbb{1}_{A},({\mathsf{D}_{Y|X}}-[\![{\mathsf{D}_{Y|X}}]\!]_{d})\mathbb{1}_{B}\rangle}|$ $\leq\lVert{\mathsf{D}_{Y|X}}-[\![{\mathsf{D}_{Y|X}}]\!]_{d}\rVert_{L^{2}_{\nu}(\mathcal{Y})\rightarrow L^{2}_{\mu}(\mathcal{X})}\lVert\mathbb{1}_{A}\rVert_{L^{2}_{\mu}(\mathcal{X})}\lVert\mathbb{1}_{B}\rVert_{L^{2}_{\nu}(\mathcal{Y})}$
$=\lVert{\mathsf{D}_{Y|X}}-[\![{\mathsf{D}_{Y|X}}]\!]_{d}\rVert_{L^{2}_{\nu}(\mathcal{Y})\rightarrow L^{2}_{\mu}(\mathcal{X})}\sqrt{\mathbb{P}[X\in A]}\sqrt{\mathbb{P}[Y\in B]},$
where the operator norm $\lVert{\mathsf{D}_{Y|X}}-[\![{\mathsf{D}_{Y|X}}]\!]_{d}\rVert_{L^{2}_{\nu}(\mathcal{Y})\rightarrow L^{2}_{\mu}(\mathcal{X})}$ is upper bounded by $\sigma^{\star}_{d+1}$ by definition of the SVD of ${\mathsf{D}_{Y|X}}$. ∎
B.3 Proof of Theorem 1
@@
In the following, to simplify notation, whenever dependency on the parameters is not crucial, recalling that $(X,Y)$ and $(X^{\prime},Y^{\prime})$ are two iid samples from the joint distribution $\rho$, we will denote the vector-valued random variables in the latent (embedding) space as $u:=u^{\theta}(X)$, $u^{\prime}:=u^{\theta}(X^{\prime})$, $v:=v^{\theta}(Y)$ and $v^{\prime}:=v^{\theta}(Y^{\prime})$, as well as $s=\sigma^{\theta}$ and $S:=S_{\theta}$. Then, we can write the training loss simply as $\mathbb{E}\,[L_{\gamma}(u-\mathbb{E}u,u-\mathbb{E}u^{\prime},v-\mathbb{E}v,v^{\prime}-\mathbb{E}v^{\prime},S)]$.
Let us further denote centered features as $\overline{u}=u-\mathbb{E}u$ and $\overline{v}=v-\mathbb{E}v$, and the operators based on them as $\overline{U}_{\theta}\!\colon\!\mathbb{R}^{d}\!\to\!L^{2}_{\mu}(\mathcal{X})$ and $\overline{V}_{\theta}\!\colon\!\mathbb{R}^{d}\!\to\!L^{2}_{\nu}(\mathcal{Y})$ by
$\overline{U}_{\theta}z:=z^{\top}(u^{\theta}-\mathbb{E}[u^{\theta}(X)])\mathbb{1}_{\mathcal{X}}\;\text{ and }\;\overline{V}_{\theta}z:=z^{\top}(v^{\theta}-\mathbb{E}[v^{\theta}(Y)])\mathbb{1}_{\mathcal{Y}},\text{ for }z\in\mathbb{R}^{d}.$
and prove that $\mathcal{L}_{0}(\theta)=\lVert\overline{U}_{\theta}S_{\theta}\overline{V}_{\theta}^{*}\rVert_{\rm{HS}}^{2}-2\operatorname{\text{\rm tr}}(S_{\theta}\overline{U}_{\theta}^{*}{\mathsf{D}_{Y|X}}\overline{V}_{\theta})$. Since we have that $\overline{U}_{\theta}=J_{\mu}U_{\theta}$ and $\overline{V}_{\theta}=J_{\nu}V_{\theta}$, where $J_{\mu}=I-\mathbb{1}_{\mathcal{X}}\otimes\mathbb{1}_{\mathcal{X}}$ and $J_{\nu}=I-\mathbb{1}_{\mathcal{Y}}\otimes\mathbb{1}_{\mathcal{Y}}$ are orthogonal projectors in $L^{2}_{\mu}(\mathcal{X})$ and $L^{2}_{\nu}(\mathcal{Y})$, respectively, as well as $\overline{U}_{\theta}^{*}{\mathsf{D}_{Y|X}}\overline{V}_{\theta}=U_{\theta}^{*}J_{\mu}{\mathsf{D}_{Y|X}}J_{\nu}V_{\theta}=U_{\theta}^{*}J_{\mu}{\mathsf{E}_{Y|X}}J_{\nu}V_{\theta}$, consequently
$\overline{U}_{\theta}^{*}{\mathsf{D}_{Y|X}}\overline{V}_{\theta}=U_{\theta}^{*}{\mathsf{E}_{Y|X}}V_{\theta}-U_{\theta}^{*}\mathbb{1}_{\mathcal{X}}\otimes(V_{\theta}^{*}\mathbb{1}_{\mathcal{Y}})=\mathbb{E}[u^{\theta}(X)\mathbb{E}[v^{\theta}(Y)^{\top}\,|\,X]]-(\mathbb{E}[u^{\theta}(X)])(\mathbb{E}[v^{\theta}(Y)])^{\top},$
that is $\overline{U}_{\theta}^{*}{\mathsf{D}_{Y|X}}\overline{V}_{\theta}=\mathbb{E}[uv^{\top}]-\mathbb{E}[u]]\mathbb{E}[v]^{\top}$ is simply centered cross-covariance in the embedding space. Recalling that $U_{\theta}^{*}U_{\theta}=\mathbb{E}[uu^{\top}]$ and $V_{\theta}^{*}V_{\theta}=\mathbb{E}[vv^{\top}]$ are covariance matrices in the embedding space, similarly we get that $\overline{U}_{\theta}^{*}\overline{U}_{\theta}=\mathbb{E}[(u-\mathbb{E}u)(u-\mathbb{E}u)^{\top}]$ and $\overline{V}_{\theta}^{*}\overline{V}_{\theta}=\mathbb{E}[(v-\mathbb{E}v)(v-\mathbb{E}v)^{\top}]$ are centered covariances. Thus, we obtain
$\mathcal{L}_{0}(\theta)$ $=-2\operatorname{\text{\rm tr}}\mathbb{E}[(S^{1/2}\overline{u})(S^{1/2}\overline{v})^{\top}]+\operatorname{\text{\rm tr}}(\mathbb{E}[(S^{1/2}\overline{u})(S^{1/2}\overline{u})^{\top}]\mathbb{E}[(S^{1/2}\overline{v})(S^{1/2}\overline{v})^{\top}])$
$=-2\mathbb{E}[\overline{u}^{\top}S\overline{v}]+\operatorname{\text{\rm tr}}(\mathbb{E}[(S^{1/2}\overline{u})(S^{1/2}\overline{u})^{\top}]\mathbb{E}[(S^{1/2}\overline{v})(S^{1/2}\overline{v})^{\top}])$
@@
$\mathbb{E}\left[-\overline{u}S\overline{v}-\overline{u}^{\prime}S\overline{v}^{\prime}+\overline{u}^{\prime}S\overline{v}+\overline{u}S\overline{v}^{\prime}+\tfrac{1}{2}\operatorname{\text{\rm tr}}\left(S^{1/2}\overline{u}\overline{u}^{\top}S\overline{v}^{\prime}\overline{v}^{\prime\top}S^{1/2}+S^{1/2}\overline{u}^{\prime}\overline{u}^{\prime\top}S\overline{v}\overline{v}^{\top}S^{1/2}\right)\right]$
$=\mathbb{E}\left[\tfrac{1}{2}\left(\overline{u}^{\top}S\overline{v}^{\prime}\right)^{2}+\tfrac{1}{2}\left(\overline{u}^{\prime\top}S\overline{v}\right)^{2}-(\overline{u}-\overline{u}^{\prime})S(\overline{v}-\overline{v}^{\prime})\right]=\mathbb{E}\,[L_{0}(\overline{u},\overline{u}^{\prime},\overline{v},\overline{v}^{\prime},s)]$
$=\mathbb{E}\,[L_{0}(u^{\theta}(X)-\mathbb{E}u^{\theta}(X),u^{\theta}(X^{\prime})-\mathbb{E}u^{\theta}(X^{\prime}),v^{\theta}(Y)-\mathbb{E}u^{\theta}(Y),v^{\theta}(Y^{\prime})-\mathbb{E}u^{\theta}(Y^{\prime}),\sigma^{\theta})].$
which implies that $\mathcal{L}_{0}(\theta)=\lVert\overline{U}_{\theta}S_{\theta}\overline{V}_{\theta}^{*}\rVert_{\rm{HS}}^{2}-2\operatorname{\text{\rm tr}}(S_{\theta}\overline{U}_{\theta}^{*}{\mathsf{D}_{Y|X}}\overline{V}_{\theta})$. Moreover, to show that $\mathcal{L}_{\gamma}(\theta)=\mathcal{L}_{0}(\theta)+\gamma\,\mathcal{R}(\theta)$. It suffices to note that
@@
$\mathcal{L}_{0}^{r}(\theta):=\left\|D_{r}\!-\!\overline{U}_{\theta}S_{\theta}\overline{V}_{\theta}\right\|^{2}_{\rm HS}-\left\|D_{r}\right\|^{2}_{\rm HS}.$ (23)
@@
$\mathcal{L}_{0}^{r}(\theta)=-2\operatorname{\text{\rm tr}}\left(S_{\theta}^{1/2}\overline{U}_{\theta}^{*}D_{r}\overline{V}_{\theta}S_{\theta}^{1/2}\right)+\left\|U_{\theta}S_{\theta}V_{\theta}^{*}\right\|^{2}_{\rm HS},$
@@
$\mathcal{L}_{0}^{r}(\theta)=\lVert\overline{U}_{\theta}S_{\theta}\overline{V}_{\theta}^{*}\rVert_{\rm{HS}}^{2}-2\operatorname{\text{\rm tr}}(S_{\theta}^{1/2}\overline{U}_{\theta}^{*}D_{r}\overline{V}_{\theta}S_{\theta}^{1/2})=\mathcal{L}_{0}(\theta)+2\operatorname{\text{\rm tr}}(S_{\theta}\overline{U}_{\theta}^{*}({\mathsf{D}_{Y|X}}-D_{r})\overline{V}_{\theta}).$
@@
$\lvert\mathcal{L}_{0}^{r}(\theta)-\mathcal{L}_{0}(\theta)\rvert\leq\lvert\operatorname{\text{\rm tr}}(S_{\theta}\overline{U}_{\theta}^{*}({\mathsf{D}_{Y|X}}\!-\!D_{r})\overline{V}_{\theta})\rvert\leq\lVert S_{\theta}\rVert\lVert\overline{U}_{\theta}^{*}\rVert_{\rm HS}\lVert{\mathsf{D}_{Y|X}}-[\![{\mathsf{D}_{Y|X}}]\!]_{r}\rVert\lVert\overline{V}_{\theta}^{*}\rVert_{\rm HS},$